[Brahms-l] Shift report 20040216 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Mon Feb 16 2004 - 16:01:10 EST
------------------------------------------------------------
Date: Mon Feb 16 16:01:10 2004, Shift: 08:00-16:00
Supervisor: DO, with: PJ
Log book#15 pages 59 - 59

Shift Summary:
------------------------------------------------------------
1. Not as bad as previous shift, we had finally a store at
   ~ 1pm, still going on 

2. Changed D5 polarity to B, current is still 1050
   Data runs:

       MRS 45 B1050 tr3            FS 4 A1/1  tr2
          10392  63k                    55k
          10394 127k                   107k
          10395 144k                   100k 
          10396 still running 

Pedestals: 10389 non TPC
           10390 TPC 
           10391 TPC (before beam store, to show Per, new
                      shifter from NBI, welcome man!) 

3. No problems from our side, detectors are cool :)
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Feb 16 16:01:30 2004

This archive was generated by hypermail 2.1.8 : Mon Feb 16 2004 - 16:01:52 EST